Actions
Operations report¶
- User Cycle 2015/04
- Began on 16th Feb and running began a week later
- MOM : Paul Kyberd (& then Ryan)
- Expert startup was mostly successful, although delayed due to
- Tour for High-level Chinese Academy of Science delegation
- Loss of dipole in ISIS which delayed beam startup for several days
- Data-taking began on 23rd February
- TOF Calibation (Durga)
- Alignment (Francois)
- Extended repeat of run 7469 for emittance studies (Victoria)
- DS PSU returned on Friday 26th and is now being installed and tested by Daresbury technicians
- Zero absorber running delayed due to problems in evacuating the absorber overlapping with unavailability of experts
(Mark Tucker was on leave) - Absorber is now being evacuated and we should be ready to take zero absorber data today. Zero absorber run-time is now
compressed (return of DS to active status may help if done in time). - Installation of LiH absorber still planned to being on Monday March 7th and to continue for the entire week.
- If the DS is operational, we could do rate studies during the evening runs of next week.
- DS Alarm incident on 29/2/2016
- Evidence that procedures are not being followed from DS Alarm incident will need to be worked through
- User Cycle 2016/01
- April user cycle : April 12th - May 20th
- Decision has (I think) been made to cancel April running (are we agreed on this? If so I will tell the MOM that they are not
required for April) We need to think carefully about what may be realistically
available by May to decide on whether we run in May (and what data we take).
- Shifts for Feb/Mar run
- BLOC signup (& size of pool) is still a concern
- Have decided to offer partial shift credit for BLOCs as incentive
- Training
- Training schedule is defined
- Will impact a little on data-taking but necessary
- http://micewww.pp.rl.ac.uk/projects/operations/wiki/TrainingPlan201504
Updated by Boyd, Steven about 7 years ago ยท 5 revisions